OpenStudy (anonymous):

How would you integarte this:

OpenStudy (anonymous):

\[\int\limits dz /(r ^{2}+z ^{2})^{3/2}\]

OpenStudy (experimentx):

is z complex no?

OpenStudy (anonymous):

no

OpenStudy (experimentx):

try usual trick ... trig substitution ... z = r tan theta

OpenStudy (australopithecus):

r^(2) is just a constant remember try to get 1/(r2+z2)^3/2 you need it in the form 1/(1 + z^(2))^(1/2) then you can use trigonometric substitution to solve so lets factor \[\frac{1}{(r^{6})^{1/2}(1+z^{6})^{1/2}}\]

OpenStudy (experimentx):

do you want it as series?

OpenStudy (australopithecus):

crud

OpenStudy (anonymous):

no

OpenStudy (australopithecus):

it should be \[\frac{1}{(r^{6})^{1/2}(1 + \frac{z^{6}}{r^{2}})^{1/2}}\]

OpenStudy (experimentx):

this doesn't look good ... from other method

OpenStudy (australopithecus):

so you have the integral: \[\frac{1}{(r^{6})^{1/2}}\int\limits_{}^{}\frac{1}{(1 + \frac{z^{6}}{r^{2}})^{1/2}}\] now use trigonometric substation to solve: here is a video that shows you how to do trig sub: http://www.youtube.com/watch?v=n4EK92CSuBE this is the list of the three methods and what trig identity to use http://en.wikipedia.org/wiki/Trigonometric_substitution

OpenStudy (anonymous):

can you do it in a *tan t method? what do you get for x^2+a^2?

OpenStudy (australopithecus):

in your case you would use tan(x)

OpenStudy (experimentx):

 let me verify